OpenStudy (anonymous):

What is the slope of the line passing through the points (–1, 3) and (4, –7) Choosed 3 over 4

OpenStudy (anonymous):

or 3 - 4

ganeshie8 (ganeshie8):

slope of a line passing thru \(\color{red}{(x_1, y_1)}\) and \( \color{green}{(x_2, y_2)}\) = \(\large \frac{\color{green}{y_2}-\color{red}{y_1}}{\color{green}{x_2}-\color{red}{x_1}}\)

OpenStudy (anonymous):

o _ O yesh

OpenStudy (anonymous):

how?? did.....

ganeshie8 (ganeshie8):

slope of a line passing thru \(\color{red}{(-1, 3)}\) and \( \color{green}{(4, -7)}\) = \(\large \frac{\color{green}{-7}-\color{red}{3}}{\color{green}{4}-\color{red}{-1}}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

its been a while since i simplified one of those

ganeshie8 (ganeshie8):

ttryyy :)

OpenStudy (anonymous):

actuality i for got how to o . O cuz my teacher in 1st grade never showed me how to :/

ganeshie8 (ganeshie8):

its okay, ima show u

OpenStudy (anonymous):

kk ;)

ganeshie8 (ganeshie8):

we want to simplify : \(\large \frac{\color{green}{-7}-\color{red}{3}}{\color{green}{4}-\color{red}{-1}}\)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

look at top, -7-3 = ?

ganeshie8 (ganeshie8):

u knw 7+3 = 10 right ? and u also knw 7-3 = 4 right ?

OpenStudy (anonymous):

(-7)-3= (-7)+(-3)= 10

ganeshie8 (ganeshie8):

Excellent ! simplify bottom also :)

OpenStudy (anonymous):

(-1)-3=(-1)+(-3)=4

ganeshie8 (ganeshie8):

\(\large \frac{\color{green}{-7}-\color{red}{3}}{\color{green}{4}-\color{red}{-1}}\) \(\large \frac{-10}{\color{green}{4}-\color{red}{-1}}\)

ganeshie8 (ganeshie8):

careful, bottom u have 4--1

OpenStudy (anonymous):

-10/4--1 is the 1 negative?

OpenStudy (anonymous):

and how do you put −104−−1

ganeshie8 (ganeshie8):

4--1 = 4 + 1 = 5

ganeshie8 (ganeshie8):

cuz, \(-\times- = +\)

OpenStudy (anonymous):

so its -10 x 5=+50?

ganeshie8 (ganeshie8):

nopes :)

OpenStudy (anonymous):

hmmm

ganeshie8 (ganeshie8):

\(\large \frac{\color{green}{-7}-\color{red}{3}}{\color{green}{4}-\color{red}{-1}} \) \(\large \frac{-10}{5} \) \(\large -2 \)

OpenStudy (anonymous):

question how did you get -2 did you dvide like so 

ganeshie8 (ganeshie8):

yes :)

ganeshie8 (ganeshie8):